# Advancements in Functorial Homological Mirror Symmetry

### v3

@@ A CohFT, then consists of a system
In the context of HMS, the WDVV equations are particularly useful for calculations involving the A-side, as explained in what follows. To define Gromov-Witten invariants of a symplectic manifold, ${\cal M}$, consider moduli spaces of stable homolorphic maps, Pan ,
¯$\overline{\cal M}_{{}_{g,n,\beta}}(M)\ \overset{def.}{=}\ \left\{u:\Sigma_{{}_{g}}\rightarrow X\ \text{holomorphic},\ u[\Sigma_{{}_{g}}]=\beta\right\}\ /\sim,$ (78)
@@
¯$\overline{\cal M}_{{}_{g,n}}\ \longleftarrow\ \overline{\cal M}_{{}_{g,n,\beta}}\ \longrightarrow\ M^{{}^{n}}$ (79)
@@ 5.1 WDVV equations
$\phi_{{}_{\alpha}}\bullet\phi_{{}_{\beta}},\phi_{{}_{\gamma}}\right>\ \overset{def.}{=}\ F_{{}_{\alpha\beta\gamma}}(t)$ (88)
@@ A Theorem asserts that given a CSR, its kernel are Lagrangian submanifolds
=${\cal{L}}\ \overset{def.}{=}\ \pi^{{}^{-1}}(0)\ =\ \bigcup_{{}_{\alpha}}\ \cal{L}_{\alpha}.$ (108)
@@
Most importantly, (114) is a statement of Langlands duality, where $K$ denotes a Kostant slice. In (114), the first equivalence is taken by Moore and Tachikawa as an axiom needed for constructing the 2D TFT in question. However, the most important application of the Moore-Tachikawa varieties consists in the case where the target of the functor (109) is $\ddot{\text{a}}$hler quotient, which is indeed the case when attempting to describe Higgs and Coulomb branches with this formalism.
@@
The starting point for explaining this is an observation originally put forward by C. Teleman, stating that, in presence of a Hamiltonian group action on the symplectic side, the HMS dual is trivial. Importantly, this does not mean that HMS is wrong. But, rather, that it requires a suitable generalisation, comprising this case. In attempting to do so, he developed fascinating works in the context of Coulomb ranches of 3D ${\cal N}=4$ SCFTs, in turn inspired by Braverman, Finkelber and Nakajima’s mathematical formulation, BFN . As explained in Section 6.1, the HMS dual of the Hamiltonian group action can be nontrivially described in terms of the Rozansky-Witten theory.
@@ 5.3 The role of Coulomb Branches for classifying 2D TFTs
6 Main techniques
The present section outlines some main techniques for determining possible extensions of the Fukaya categories that are of particular relevance for String Theory setups. It is structured as follows:
@@ 6 Main techniques
    Section 6.1 outlines the abelianisation technique, one of the main techniques in addressing HMS for singular varieties. Concretely relying on the quantum mechanics of the cohomology of Rozansky-Witten theory on the B-side. Under ordinary HMS correspondence, abelianisation induces additional Lagrangian submanifolds on the A-side.
@@
    Section 6.2 deals with Koszul duality [^13] Originally introduced by Kontsevich as a generalisation of Lie theory. applied to the setup of Higgs and Coulomb branches.
  3. 3.
    Section 6.3 is mostly an overview of part of Sheridan’s work, She , in determining the HMS dictionary of the pair of pants, and the need to account for Lagrangian immersions.
@@ 6.1 Abelianisation
6.2 Koszul duality
@@ 6.3 Lagrangian immersion
$L^{{}^{n}}:S^{{}^{n}}\ \rightarrow\ {\cal P}^{{}^{n}}\ \ \ \ ,\ \ \ \ \ {\cal P}^{{}^{n}}\ =\ \left\{\sum_{{}_{j=1}}^{{}^{n+2}}z_{{}_{j}}=0\right\}\ \subset\ \mathbb{C}^{{}^{n+2}}\backslash\ \bigcup_{{}_{j}}\ \left\{z_{{}_{j}}=0\right\}.$ (127)
@@ A Theorem in She states that there exists a fully-faithful $A_{{}_{\infty}}$-embedding
$CF(L,L)\ =\ \text{Span}\left\{1,X,Y,Z,\overline{X},\overline{Y},\overline{Z},\text{pt}\right\}.$ (132)
@@
Denoting by ${\cal L}$ the collection of lifts of the Seidel Lagrangian $\overline{\mathbb{L}}$ in a manifold cover, there is an isomorphism
@@
then ${\cal L}$ split-generates $Fuk(M)$.
7 Conclusions and Outlook
The essential theme this article is the importance of further generalising the mathematical structure underlying functorial HMS, mostly motivated from its applicability to String Theory, with quiver gauge theories playing a crucial role in the setting. However, as we have seen, generalising this construction also constitutes a promising setting for pure mathematics advancements in their own right.
The present work was structured as follows: Section 2 introduced symmetric obstruction theories as shifted symplectic structures, DO ; PTVV . Section 3 briefly overviewed a recent work by Katzarkov, Kontsevich and Sheshmani, KKS , pertaining the categorification of DT-invariants for cases involving non-smooth moduli spaces, relative CY structures and spherical functors. In Section 4 we then turned to the topological string theory formulation of Section 3, mostly related to Pan ; AOSV . Section 5 applied the tools outlined in the previous sections to a particular setup of interest for its applicability to String Theory, namely that of Moore-Tachikawa varieties, MT ; MS ; P ; CM . Particular emphasis was placed on the role of Coulomb branches for classifying 2D TFTs, XY ; CT ; CT1 ; BFN . In Section 6 we then turned to some of the main techniques used for extending HMS for singular varieties.
Along the treatment, several possible directions for further development have been identified. We will be reporting about any advancement in this regard in due course.
Acknowledgements
I wish to thank Gregory Moore, Cumrun Vafa, and Edward Witten for insightful discussions and questions raised at different stages of the present work. Early stages of it have been presented at the Chern Mathematics Institute (Tianjin), KITS Beijing, Fudan University Maths Department (Shanghai), and NYU in Abu Dhabi in occasion of Strings 2025. An almost ultimated version was presented as a plenary talk at the Geometry and Physics Workshop in January 2025 at TSIMF, Sanya, organised by Professor Shing-Tung Yau and Professor Bong Lian. I also wish to acknowledge hospitality from IMSA, Miami, in occasion of the Homological Mirror Symmetry Conference in 2025 during the concluding stage of this work. My research is fully-funded by SIMIS.
Appendix A Essential Concise Background
  1. 1.
  2. 2.
  3. 3.
  4. 4.
A.1 Homological Mirror Symmetry (HMS)
$D^{{}^{b}}\text{Fuk}(X)\ \simeq\ D^{{}^{b}}\text{Coh}(\hat{X})$ (135)
A.2 Functorial HMS
${\cal F}:D^{{}^{b}}\text{Fuk}(X)\ \longrightarrow\ D^{{}^{b}}\text{Coh}(\hat{X}).$ (136)
For the case in which $X,\hat{X}$ are mirror CY manifolds, (136) is essentially stating that given a certain symplectic manifold, we can obtain $\ddot{\text{a}}$hler manifold, belonging to another CY variety. It is known that, from dimensional reduction of 10 D string theory, there is a rich web of dualities arising according to the specific properties of the manifold on which compactification is performed.
A.3 The LG/CY correspondence
@@
$d\ =\ w_{{}_{1}}+...+w_{{}_{N}}$ (137)
Two objects can be defined:
    $\mu_{{}_{d}}\ =\ \{g\in\mathbb{C}^{{}^{\times}}|g^{{}^{d}}=1\}$ (138)
    $(x_{{}_{1}},...,x_{{}_{N}})\ \mapsto\ (g^{{}^{w_{{}_{1}}}}x_{{}_{1}},...,g^{{}^{w_{{}_{N}}}}x_{{}_{N}}),$
    and a $\mu_{{}_{d}}$-invariant potential
    ${\cal W}:\ \mathbb{C}^{{}^{N}}\ \longrightarrow\ \mathbb{C}.$ (139)
    $\mathbb{P}(\underline{w})\ =\ \mathbb{P}(w_{{}_{1}},...,w_{{}_{N}}),$ (140)
    which is quasi-smooth, namely smooth as a stack.
    These two models arise from a modified GIT-quotient
A.4 MF of the pair of pants
[…]
$\mathbb{T}\ \overset{def.}{=}\ \text{Hom}\left(M,\mathbb{C}^{{}^{*}}\right).$ (142)
The B-model on $\left(\text{Spec}(R),{\cal W}\right)$ is given by the triangulated category of singularities of ${\cal W}^{{}^{-1}}(0)$, the latter being quasi-equivalent to MF$(R,{\cal W})$
[…]
=${\cal B}\ \overset{def.}{=}\ \text{Hom}^{{}^{*}}_{{}_{\text{MF}(R,{\cal W})}}\left({\cal O}_{{}_{o}},{\cal O}_{{}_{o}}\right).$ (144)
It is instructive to see how the $\mathbb{T}$-action enters the picture.
MF corresponding to ${\cal O}_{{}_{o}}$ is the Koszul resolution of ${\cal O}_{{}_{o}}$
$R\otimes\Lambda^{{}^{*}}\tilde{M}$ (145)
with the deformed differential
$\delta\ \overset{def.}{=}\ \iota_{{}_{n}}\ +\ v\wedge$ (146)
where
$u$ =$\overset{def.}{=}\ \sum_{{}_{j=1}}^{{}^{n+2}}z_{{}_{j}}\theta_{{}_{j}}^{{}^{\text{V}}}\ \in\ R\otimes\ \Lambda^{{}^{*}}\tilde{M}^{{}^{\text{V}}}$ (147)
$v$ =$\overset{def.}{=}\ \sum_{{}_{j=1}}^{{}^{n+2}}a_{{}_{j}}\frac{{\cal W}}{z_{{}_{j}}}\theta_{{}_{j}}\ \in\ R\otimes\ \Lambda^{{}^{*}}\tilde{M}.$
where
$\delta\ =\ \sum_{{}_{j=1}}^{{}^{n+2}}z_{{}_{j}}\ \left(\frac{\partial}{\theta_{{}_{j}}}+a_{{}_{j}}\frac{{\cal W}}{z_{{}_{j}}}\theta_{{}_{j}}\ \in\ R\otimes\ \Lambda^{{}^{*}}\tilde{M}\right)$ (149)
with the differential
$d\ \overset{def.}{=}\ [\delta,-]$ (152)
Appendix B Stability of Higgs bundles
A Higgs bundle is a holomorphic bundle $\left(E,\overline{\partial}_{{}_{E}}\right)$ and a holomorphic section $\phi$ of End$(E)\otimes\ {\cal K}_{{}_{X}}$, where
¯$\overline{\partial}_{{}_{E}}:\ \Gamma(X,E)\ \rightarrow\ \Gamma\left(X,E\otimes\Omega^{{}^{(0,1)}}(x)\right),$ (153)
and $\Gamma(X,E)$ denotes the space of all sections of $E$.
and stable if (154) is a strict inequality.

### v4

@@ A CohFT, then consists of a system
[…]
$\overline{\cal M}_{{}_{g,n,\beta}}(M)\ \overset{def.}{=}\ \left\{u:\Sigma_{{}_{g}}\rightarrow X\ \text{holomorphic},\ u[\Sigma_{{}_{g}}]=\beta\right\}\ /\sim,$ (105)
@@
$\overline{\cal M}_{{}_{g,n}}\ \longleftarrow\ \overline{\cal M}_{{}_{g,n,\beta}}\ \longrightarrow\ M^{{}^{n}}$ (106)
@@ 5.1 WDVV equations
$\left<\phi_{{}_{\alpha}}\bullet\phi_{{}_{\beta}},\phi_{{}_{\gamma}}\right>\ \overset{def.}{=}\ F_{{}_{\alpha\beta\gamma}}(t)$ (115)
@@ A Theorem asserts that given a CSR, its kernel are Lagrangian submanifolds
${\cal{L}}\ \overset{def.}{=}\ \pi^{{}^{-1}}(0)\ =\ \bigcup_{{}_{\alpha}}\ \cal{L}_{\alpha}.$ (135)
@@
[…]
@@
The starting point for explaining this is an observation originally put forward by C. Teleman, stating that, in presence of a Hamiltonian group action on the symplectic side, the HMS dual is trivial. Importantly, this does not mean that HMS is wrong. But, rather, that it requires a suitable generalisation, comprising this case. In attempting to do so, he developed fascinating works in the context of Coulomb ranches of 3D ${\cal N}=4$ SCFTs, in turn inspired by Braverman, Finkelberg and Nakajima’s mathematical formulation, BFN . As explained in Section 6.1, the HMS dual of the Hamiltonian group action can be nontrivially described in terms of the Rozansky-Witten theory.
@@ 5.3 The role of Coulomb Branches for classifying 2D TFTs
6 Main technique
The present section outlines a major technique for determining possible extensions of the Fukaya categories that are of particular relevance for String Theory setups.  [^15] We refer the interested reader to Appendix C and D for further techniques that have been explored in the literature. It instructive to note that all three techniques are intertwined with each other. However, we will only be focusing on one of them for the purpose of this article.This Section is structured as follows:
@@ 6 Main technique
    Section 6.1 outlines the abelianisation technique, one of the main techniques in addressing HMS for singular varieties. Concretely relying on the quantum mechanics of the cohomology of Rozansky-Witten theory on the B-side, we explain how it induces additional Lagrangian submanifolds on the A-side.
@@
    Section 6.2 explains the relation between gauging and deformation quantisation to the construction of virtual fundamental classes. In so doing, we explain why it is unsuitable for the nonabelian case, and why we should expect currently-known techniques to be further extended.
@@ 6.1 Abelianisation
6.2 Relation to the virtual fundamental class
We now turn to explaining how the abelian/nonabelian cases outlined above are related to the categorification of invariants introduced in Section 3. In order to do so, we need to set some preliminary notation, mostly taken from CT2 ; KRS . We denote by $\mathfrak{Coh}(\mathcal{X})$ the category of coherent sheaves of modules over a scheme or variety, $\mathcal{X}$.
For torus action, which is abelian, gauging the Fukaya category of such a variety, $\mathcal{F}(\mathcal{X})$, amounts to enrich it from $\mathfrak{Coh}\left(T_{{}_{\mathbb{C}}}^{{}^{\text{V}}}\right)$-module [^16] Consisting of a 2-category $\mathfrak{Coh}(\mathcal{X}^{{}^{\text{V}}})$ determined by the map $\pi:\ \mathcal{X}^{{}^{\text{V}}}\ \longrightarrow\ T_{{}_{\mathbb{C}}}^{{}^{\text{V}}}$ (149) whose objects are Lagrangian submanifolds. These are the 2D TFTs at the boundary of the RW-theory (the 3D bulk TFT). to an object in $\sqrt{\mathfrak{Coh}\ }\left(T^{{}^{*}}T_{{}_{\mathbb{C}}}^{{}^{\text{V}}}\right)$ [^17] Such objects are the KRS-branes living in the 3D TFT. Such category is the one with $T$-action that sees the germ of the object near the 0-section (drawn in orange in the picture on the left-hand-side of Figure 10.. In the nonabelian case, though, gauging brings a holomorphic algebraic manifold which is not a cotangent bundle [^18] The reason for it not being a cotangent bundle as a whole is that it involves two projections.
$T^{{}^{*}}G_{{}_{\mathbb{C}}}^{{}^{\text{V}}}\ \equiv\ BFM\left(G^{{}^{\text{V}}}\right)\ \longleftarrow\ BFM\left(G^{{}^{\text{V}}}\right)\ \underset{\mathfrak{t}_{{}_{\mathbb{C}}}/W}{\times}\ \mathfrak{t}_{{}_{\mathbb{C}}}\ \longrightarrow\ T^{{}^{*}}T_{{}_{\mathbb{C}}}^{{}^{\text{V}}}$ (150)
The crucial difference between (150) and the abelian setup, is that the former does not allow for a direct morphism between $T^{{}^{*}}G_{{}_{\mathbb{C}}}^{{}^{\text{V}}}$ and $T^{{}^{*}}T_{{}_{\mathbb{C}}}^{{}^{\text{V}}}$. As explained in Section 6.2.1, this means it is not possible to define a SOT by means of any of the techniques outlined in Section 2. Indeed, the presence of two distinct projections that are not isomorphic to each other in (150) implies that, this time, the definition of the KRS-objects involves two distinct $\sqrt{\mathfrak{Coh}\ }(X)$.
6.2.1 $\sqrt{\mathfrak{Coh}\ }(X)$
A symmetric complex $(\mathbb{E},\theta,\text{or})$ on an Artin stack, $\chi$, is a perfect complex $\mathbb{E}$ of amplitude [-2,0] consisting of the following data:
  1. 1.
    A non-degenerate symmetric form, $\theta$ on $\mathbb{E}$, namely a morphism
    $\Theta:\ \theta_{{}_{\chi}}\ \rightarrow\ \left(\mathbb{E}\otimes\mathbb{E}\right)[-2]$ (151)
    in $D\text{Coh}(X)$, invariant under transport such that the induced morphism
    $\mathbb{E}^{{}^{\text{V}}}\ \rightarrow\ \mathbb{E}[-2]$ (152)
    is an isomorphism.
  2. 2.
    An orientation, or, of $\textbb{E}$, namely an isomorphism
    $\text{or}:\ \mathcal{O}_{{}_{\mathcal{X}}}\ \xrightarrow{\simeq}\ \text{det}\left(\mathbb{E}\right)$ (153)
    such that
    $\text{det}\left(\iota_{{}_{\theta}}\right)\ =\ \text{or}\circ\ \text{or}^{{}^{\text{V}}}$ (154)
For any symmetric complex $\mathbb{E}$, there is a quadratic function
$q_{{}_{\mathbb{E}}}:\ \mathfrak{C}_{{}_{\mathbb{E}}}\ \rightarrow\ \mathbb{A}_{{}_{\mathcal{X}}}^{{}^{1}}$ (155)
with $\mathfrak{C}_{{}_{\mathbb{E}}}$ denoting the virtual normal cone of $\mathbb{E}$. When $\mathbb{E}=E[1]$ for SO(,)($\mathbb{C}$)-bundle, $E$, the function $q_{{}_{\mathbb{E}}}$ is given by the quadratic form on $E$. Denoting by $\mathfrak{Q}(\mathbb{E})$ the 0-locus of the quadratic function (155). The SOT
$\phi:\ \mathbb{E}\ \rightarrow\ L_{{}_{f}}$ (156)
for the Deligne-Mumford morphism
[…]
induces an embedding
[…]
leading to the square root virtual pullback map
$\sqrt{f!\ }:\ CH_{{}_{\bullet}}(\mathcal{Y})\ \rightarrow\ CH_{{}_{\bullet}}\left(\mathfrak{C}_{{}_{f}}\right)\ \rightarrow\ CH_{{}_{\bullet}}(\mathfrak{Q}(\mathbb{E}))\ \xrightarrow{\sqrt{0^{{}^{!}}_{{}_{\mathfrak{Q}(\mathbb{E})}}\ }}\ CH_{{}_{\bullet}}(\mathcal{X})$ (159)
When $\mathcal{X}$ is a separated Deligne-Mumford stack over $\mathbb{C}$ with structure morphism
$a_{{}_{\mathcal{X}}}:\ \mathcal{X}\ \rightarrow\ \text{Spec}\ \mathbb{C}$ (160)
and an isomorphic SOT, (156), the virtual fundamental class is defined as follows
$\mathcal{X}]^{{}^{\text{vir}}}\ \overset{\text{def.}}{=}\ \sqrt{a^{{}^{!}}_{{}_{\mathcal{X}}}\ }\left[\text{Spec}\ \mathbb{C}\right]\ \in\ CH_{{}_{\text{Vdim}(\mathcal{X})}}(\mathcal{X})$ (161)
As long as we deal with the abelian setup (cf. the left-hand-side picture in Figure 10), we can always define (161), and the degeneracy formula can in turn be achieved since the two sides will be related by deformation quantisation. In such case, we can apply the shifted symplectic structure technique to DG-modules and calculate invariants without the need to constrain the analysis to ideal sheaves.
6.3 Main points
From our treatment, we are therefore led to make the following observations:
  1. 1.
    We need to understand how the virtual fundamental class can be defined in the nonabelian case, namely where two distinct $\sqrt{\mathfrak{Coh}\ }(X)$ are involved. The main issue being that the two sides are not related by a single deformation-quantisation step.
  2. 2.
    How to recover the abelianised version of 1. It is reasonable to expect that this should introduce new sheaves.
  3. 3.
    Physically, this should correspond to new fields in the corresponding String Theory setup.
  4. 4.
    Shifted symplectic structures do not provide the ultimate answer to address this question, due to the constraining requirements on the sheaves involved for the calculation of invariants. Even upon relaxing the ideal sheaf assumption for achieving the degeneracy formula, we still cannot claim anything regarding the nonabelian setup. So far, only the abelian case can be addressed with such tool. The A-side of homological mirror symmetry therefore requires further advancements.
  5. 5.
    Abelianisation is the ultimate goal, but there is clearly a formal gap to be filled, which should be filled, and that should clearly provide very promising useful insights for, both furthering the String Theory realisation of effective field theories, as well as the development of mathematical tools for enumative geometry and homological mirror symmetry.
We are currently working towards identifying the nonabelian counterpart of (161), understanding its mathematical structure, and the implications of its abelianisation to the degeracy formula as well as to its D-brane realisation. We plan to report any advancement in this regard in due course.
7 Conclusions and Outlook
The essential theme in this article is the importance of further generalising the mathematical structure underlying functorial HMS, mostly motivated from its applicability to String Theory, with quiver gauge theories playing a crucial role in the setting. However, as we have seen, generalising this construction also constitutes a promising setting for pure mathematics advancements in their own right.
Our main finding consists in having related stability and transversality in the degeneracy formula for non-ideal sheaves to nonabelian gauging in the RW-theory setup.
The present work was structured as follows: Section 2 introduced symmetric obstruction theories as shifted symplectic structures, DO ; PTVV . Section 3 briefly overviewed a recent work by Katzarkov, Kontsevich and Sheshmani, KKS , pertaining the categorification of Donaldson-Thomas (DT) invariants for cases involving non-smooth moduli spaces, relative CY structures and spherical functors. The example we outlined falls within the theme of categorification of AG. We highlighted the types of mathematical structures we needed to apply to achieve the degeneracy formula, and how it relates to deformation quantisation. In Section 4 we then turned to the topological string theory formulation of Section 3, mostly related to Pan ; AOSV . Here we stressed the role played by different brane types in the braneweb for which the topological partition function is being calculated, and how they are encoded in a 2-functor corresponding to Gromov-Witten (GW) invariants [^19] DT-invariants are an equivariant formulation of GW-invariants. Calculations coincide for CY3-folds.. Cobordism calculations, though, highlighted that such invariants restrict to critical points of a topological Morse function. More interesting configurations should be expected to arise when attempting to analyse geometric structures arising at arbitrary points of the flow. In Section 5, we applied the tools outlined in the previous sections to a specific setup of interest for its applicability to String Theory, namely that of Moore-Tachikawa varieties, MT ; MS ; P ; CM . Particular emphasis was placed on the role of Coulomb branches for classifying 2D Topological Field Theories (TFTs), XY ; CT ; CT1 ; BFN . The reason for doing so is that such 2D TFTs constitute the boundary of a 3D TFT described by Rozansky and Witten for describing the category of branes associated to the B-model and their deformation. Section 6 turned to a major technique used for extending HMS for singular varieties, namely that of abelianisation. In so doing, we explained the relation between gauging and deformation quantisation, (1), and the difference between the abelian and the nonabelian setup, highlighting how this relates to the definition of the virtual fundamental class in the degeneracy formula. We concluded by outlining some main open questions and currently ongoing work by the same author towards extending the formalism in question. We plan to report any advancement in this regard in due course.
Acknowledgements
The author wishes to thank Gregory Moore, Cumrun Vafa, and Edward Witten for insightful discussions and questions raised at different stages of the present work. Early stages of it have been presented at the Chern Mathematics Institute (Tianjin), KITS Beijing, Fudan University Maths Department (Shanghai), and NYU in Abu Dhabi in occasion of Strings 2025. An almost ultimated version was presented as a plenary talk at the Geometry and Physics Workshop in January 2025 at TSIMF, Sanya, organised by Professor Shing-Tung Yau and Professor Bong Lian. Its completed version was first presented at a Joint Tsinghua-BIMSA Seminar on Symplectic Geometry in June 2025. The author also acknowledges hospitality from IMSA, Miami, in occasion of the Homological Mirror Symmetry Conference in 2025 during the concluding stage of this work. Last but not least, the author wishes to thank the anonymous referee, for insightful comments and suggestions, that helped towards improving this work towards its present form. The author’s research is funded by SIMIS.
Appendix A Essential Concise Background
  1. 1.
  2. 2.
  3. 3.
  4. 4.
A.1 Homological Mirror Symmetry (HMS)
[…]
A.2 Functorial HMS
[…]
A.3 The LG/CY correspondence
[…]
$d\ =\ w_{{}_{1}}+...+w_{{}_{N}}$ (164)
Two objects can be defined:
    $\mu_{{}_{d}}\ =\ \{g\in\mathbb{C}^{{}^{\times}}|g^{{}^{d}}=1\}$ (165)
    $(x_{{}_{1}},.,x_{{}_{N}})\ \mapsto\ (g^{{}^{w_{{}_{1}}}}x_{{}_{1}},.,g^{{}^{w_{{}_{N}}}}x_{{}_{N}}),$
    and a $\mu_{{}_{d}}$-invariant potential
    ${\cal W}:\ \mathbb{C}^{{}^{N}}\ \longrightarrow\ \mathbb{C}.$ (166)
    $\mathbb{P}(\underline{w})\ =\ \mathbb{P}(w_{{}_{1}},...,w_{{}_{N}}),$ (167)
    which is quasi-smooth, namely smooth as a stack.
    These two models arise from a modified GIT-quotient
A.4 MF of the pair of pants
${\cal P}^{{}^{n}}\ \ \ \overset{\text{HMS}}{<-------->}\ \ \ \left(\text{Spec}(R),{\cal W}\right),$ (168)
$\mathbb{T}\ \overset{def.}{=}\ \text{Hom}\left(M,\mathbb{C}^{{}^{*}}\right).$ (169)
[…]
$L^{{}^{n}}\ \ \ \overset{\text{HMS}}{<-------->}\ \ \ {\cal O}_{{}_{o}}\ \in\ D_{{}_{sing}}^{{}^{b}}\left({\cal W}^{{}^{-1}}(0)\right).$ (170)
${\cal B}\ \overset{def.}{=}\ \text{Hom}^{{}^{*}}_{{}_{\text{MF}(R,{\cal W})}}\left({\cal O}_{{}_{o}},{\cal O}_{{}_{o}}\right).$ (171)
It is instructive to see how the $\mathbb{T}$-action enters the picture.
MF corresponding to ${\cal O}_{{}_{o}}$ is the Koszul resolution of ${\cal O}_{{}_{o}}$
$R\otimes\Lambda^{{}^{*}}\tilde{M}$ (172)
with the deformed differential
$\delta\ \overset{def.}{=}\ \iota_{{}_{n}}\ +\ v\wedge$ (173)
where
$u$ $\overset{def.}{=}\ \sum_{{}_{j=1}}^{{}^{n+2}}z_{{}_{j}}\theta_{{}_{j}}^{{}^{\text{V}}}\ \in\ R\otimes\ \Lambda^{{}^{*}}\tilde{M}^{{}^{\text{V}}}$ (174)
$v$ $\overset{def.}{=}\ \sum_{{}_{j=1}}^{{}^{n+2}}a_{{}_{j}}\frac{{\cal W}}{z_{{}_{j}}}\theta_{{}_{j}}\ \in\ R\otimes\ \Lambda^{{}^{*}}\tilde{M}$.
where
[…]
with the differential
$d\ \overset{def.}{=}\ [\delta,-]$ (179)
Appendix B Stability of Higgs bundles
[…]
$\overline{\partial}_{{}_{E}}:\ \Gamma(X,E)\ \rightarrow\ \Gamma\left(X,E\otimes\Omega^{{}^{(0,1)}}(x)\right),$ (180)
and $\Gamma(X,E)$ denotes the space of all sections of $E$.
and stable if (181) is a strict inequality.
Appendix C Koszul duality
@@ Appendix D Lagrangian immersion
[…]
@@ A Theorem in She states that there exists a fully-faithful $A_{{}_{\infty}}$-embedding
[…]
@@
[…]